5. Jane has twelve stamps worth 25 cents each and eight stamps worth 13 cents each? She selects five stamps without looking. What is the probability that she takes three of the 25 cent stamps and two of the 13 cent stamps?

Answer by ewatrrr(24785)

 
Hi
Jane has 20 stamps: twelve stamps worth 25 cents each and eight stamps worth 13 cents each?
Selects 5
P(3(25¢) and 2(13¢) = 12C3*8C2/20C5 = 220*28/15504 = .3973